# In-Situ Manipulation of Superconducting Properties via Ultrasonic Excitation

### v1

@@ I.6 Conclusion
This study demonstrates a novel method for manipulating both orbital and lattice excitations (phonons) in materials. This is achieved through the application of ultrasonic excitation, which allows for in-situ control of these degrees of freedom. The effectiveness of this technique was confirmed by observing the influence of ultrasonic excitation on the magnetic properties of a paramagnetic salt, Gd2O3. This material was chosen due to its lack of orbital angular momentum (L=0), ensuring that any observed changes are solely attributable to phonon modulation. The absence of local heating effects due to ultrasonic excitation was confirmed by the unchanged magnetic susceptibility at high excitation levels.
Furthermore, we demonstrated the ability to manipulate the energy gap of a conventional superconductor (Nb) using ultrasonic excitation. This finding holds potential for applications in various superconducting quantum devices. This work also opens new avenues for investigating the role of phonons in high-Tc cuprate superconductors. Specifically, it suggests that the observed decrease in the superconducting transition temperature of cuprates under ultrasonic excitation may be linked to the manipulation of the antiferromagnetic network by phonons.
I.7 Acknowledgment
@@
This paper is dedicated to the memory of the late Dr. Alok Banerjee. The authors would also like to acknowledge Dr. Santanu De for his help with measurements and Er. Kranti Kumar Sharma for insightful discussions.

### v2

@@ Abstract
This study establishes ultrasonic excitation as a novel and effective method for manipulating both orbital and lattice excitations (phonons) in materials, enabling in-situ control of these crucial degrees of freedom. The impact of this technique was confirmed through observations of phonon-driven changes in the magnetic properties of Gd2O3, a paramagnetic salt chosen for its lack of orbital angular momentum. Crucially, the absence of local heating effects from ultrasonic excitation was verified, ensuring the observed changes were intrinsic to phonon modulation.
Furthermore, we successfully demonstrated the manipulation of the energy gap in a conventional superconductor (Nb) using this method, opening exciting possibilities for applications in superconducting quantum devices. Our findings also pave the way for new investigations into the role of phonons in high-Tc cuprate superconductors, suggesting a link between ultrasonic excitation, phonon modulation of the antiferromagnetic network, and the observed reduction in superconducting transition temperature.This technique offers a compelling alternative to conventional methods for manipulating critical thermodynamic parameters in superconductors, with advantages in controllability and time efficiency. Potential applications include decoupling superconductivity from two-level systems (TLS) in qubit-based experiments, replacing persistent switch heaters in superconducting magnets, and other scenarios requiring precise and dynamic tuning of superconducting properties. This work highlights the broad utility of ultrasonic excitation as a tool for manipulating and tuning material properties, with significant implications for both fundamental research and technological applications.
@@
This paper is dedicated to the memory of the late Dr. Alok Banerjee. The authors would also like to acknowledge Er. Kranti Kumar Sharma for insightful discussions.